Adeel age is 67 years and his son Hameed is 29,in how many years will Hameed be exactly half his father age?
A) 7
B) 4
C) 9
D) 5

Answer by mananth(16946)

after x years he will be half the age
1/2(67+x) = (29+x)
67+x=2(29+x)
67+x=58+2x
67-58 =2x-x
9=x
After 9 years